OpenStudy (mellani):

Find the average rate of change for the given function from x = −2 to x = 1.

OpenStudy (mhchen):

You would need to find the output of f(1) and f(-2) then set it up like this: \[\frac{f(x_{2})-f(x_{1})}{x_{2}-x_{1}}=\frac{f(1)-f(-2)}{1-(-2)}\]

OpenStudy (mathmale):

Please share with us the function in question: f(x) = ???